20% of people are vegetarians. 70% of vegetarians like broccoli. 20% of vegetarians like

carrot. 10% of vegetarians like cucumber. 40% of meat eaters like broccoli. 60% of meat

eaters like carrot. What is the probability that a man who likes carrot is a meat eater?


1
Expert's answer
2021-09-28T01:57:46-0400

To answer the question we will find the amount of people that like carrot, the amount of people that eat meat and like carrot, and then find the probability.

Let X-amount of people. Then, there are:

1) "a=0.2*0.2*x=0.04x" people that are vegeterians and like carrot

2) "b=0.8*0.6*x= 0.48x" people that are non-vegeterians and like carrot.

So, there are 0.52x people that like carrot and 0.48x of them are meat eaters. The answer is "0.48x\/0.52x=12\/13"